# SoK: “Interoperability vs Security” Arguments: A Technical Framework

### v1

@@ 2 Our Framework
Security implications of existing
@@
Table 1: Some Key Differences Between Security Engineering Concerns and Security Vetting Concerns
“Security vs. interoperability” arguments
Our framework’s focus is on contexts in which a company that has significant market power [^5] What counts as a relevant degree of market power for antitrust action depends on the competition laws and competition authorities in different jurisdictions. Our framework and contributions are agnostic to the precise threshold of market power prescribed by any given jurisdiction. invokes security-related concerns to argue for preventing, deterring, or making economically onerous third-party developers’ efforts to interoperate with the company’s platforms or systems. Our scope thus includes the invocation of security concerns to argue against legislation and regulation that encourages or mandates that such a company interoperate with third-party developers’ products or services. We call these “security vs. interoperability” arguments.
We have found that “security vs. interoperability” arguments are often made in public statements, including developer- or consumer-facing material. They have also been used in the context of regulatory actions and lawsuits in some instances. In this paper, we focus on the arguments made in contexts where a formal legal or regulatory action has been taken, as they are comparatively well documented, and as the potential impact in such cases is more concrete. That said, our analytical framework generalizes to contexts not featuring any formal legal or regulatory actions.
Scope of “interoperation”
When products and services produced by one company (e.g., smartphones) support or facilitate user engagement with products and services developed by a third-party, (e.g., independent apps), we say they interoperate and are interoperable.
For our purposes, interoperation comprises both technical interoperability (e.g., what are the requirements for an app to be technically capable of running on a particular platform?) and interoperation policy (e.g., which of the technically interoperable apps does a platform permit as a matter of policy?). Thus, if a third-party developer is blocked from offering their app on a platform even though it is technically compatible, we consider this an issue of interoperation (but not of specifically technical interoperability).
In many engineering contexts, technical interoperability would be the primary topic implicitly referenced when discussing interoperation. In many law and policy contexts, references to interoperability often encompass interoperation policy or interoperation more broadly. As such, the broader scope is necessary to capture the full range of concerns that feature in ongoing “security vs. interoperability” discourse.
Other terms
We write company or platform to refer to the developer of the main platform and third party to refer to those developers looking to interoperate. For brevity, we may write products to encompass both products and services.
2.1 Our Taxonomy
We now introduce our main taxonomy, which distinguishes two broad types of “security vs. interoperability” arguments as raising either security engineering concerns (§2.1.1) or security vetting concerns (§2.1.2), summarized in Table 1, alongside a third hybrid category that raises concerns incorporating aspects both of the earlier two categories (§2.1.3).
Our taxonomy clarifies the different types of security concerns that may arise in response to the prospect of introducing interoperation between products or systems which did not previously interoperate. The technical, economic, and policy context in which these concerns are raised is critical to our categorization and analytical framework. In particular, the status quo of how the systems in question work before interoperation is introduced, the nature of the technical security guarantees the systems already provide, and the security policies and trade-offs the companies already have in place, are essential considerations.
2.1.1 Security Engineering Concerns
When considering interoperating systems that are currently not technically interoperable, there are contexts in which preserving security alongside interoperation requires that some aspect of the company’s product be non-trivially changed. How to change the product to preserve the security guarantees currently offered by the company’s product, while also supporting interoperation, may or may not be clear.
@@ 2.1.1 Security Engineering Concerns
Template of a Security Engineering Concern From a platform’s perspective 1. Interoperating would require a building new technical mechanism or architecture for our system to interface with other systems. 2. Our system has existing technical security guarantees. Designing a new architecture to preserve these guarantees is an engineering task that is costly in time, effort, and coordination with others—and may require innovation, so success is not certain. 3. Therefore, interoperating secure systems is more costly than interoperating systems in general, and if we don’t get the engineering right, the security of billions of users could be undermined.
This template argument is logically coherent and highlights legitimate security concerns (as will the other templates that follow). It can thus be tempting, as a security expert asked to evaluate this argument, to pronounce it sound and move on.
But our focus in this paper is not on the logical soundness of these arguments or whether they highlight legitimate security issues; they usually do both of those things. [^6] Of course, “security vs. interoperability” arguments that did not meet these criteria would also be concerning and possibly misleading. We consider them out of scope not only because they are less prevalent, but also because they can be analyzed and debunked from a purely technical perspective. Instead, we focus on the broader implications of how this type of argument can be and is in fact used in “security vs. interoperability” contexts. As we will see throughout the paper, analyzing this broader context illuminates how the different types of arguments in our taxonomy can be used to promote security and/or to push for other goals unrelated to security.
2.1.2 Security Vetting Concerns
Security vetting concerns arise in contexts where there is already some level of interoperation supported by a system’s technical architecture. Many devices and platforms support an ecosystem of apps, software, and connected devices that are integral parts of their functionality, but that are developed by third parties. Examples include platforms, browsers, or operating systems on which third-party software can run, and platforms that offer API access to third-party developers.
Once such an architecture is in place, there remains the question of which third parties a platform is willing to interoperate with. Downloads and other connections can open a vector of attack for malware, spyware, and other undesirable outcomes for both the user and the device. Companies may vet and control which apps, systems, or software are permitted to interact with their devices with a view to preventing or mitigating such undesirable security outcomes (although these decisions might be motivated by other goals too).
@@ 2.1.2 Security Vetting Concerns
Template of a Security Vetting Concern From a platform’s perspective 1. Interoperating would mean our users and systems interacting with third parties and third-party software. 2. Third parties and third-party software may make different decisions about security than us. We cannot guarantee that third-party software would adhere to the same standards as software we have developed or carefully examined ourselves. 3. Therefore, allowing third-party software on our platform could undermine billions of users’ security.
@@
The above template is so generic that it could apply to any interoperation. As such, we find that vetting arguments often boil down to arguments directly against interoperation.
@@
Vetting third parties can be an important way to protect users, but also concentrates power and discretion of incumbent companies over ecosystems they control. Companies in this position can use their role as security guarantor to charge developers for access or prevent competitors from reaching the market. Their interoperation policy choices strike a tradeoff between security, interoperation, and other business considerations that can be opaque, discretionary and (necessarily) not amenable to precise technical specification.
@@
Due to the discretionary tradeoffs involved, arguments that security considerations necessitate a particular form of vetting or security measures generally cannot be substantiated by purely technical reasoning; rather, they must refer to existing interoperation policies and the discretionary choices already present therein. This contrasts with security engineering concerns, which technical reasoning may suffice to substantiate.
@@
Let us concretize the last point with a few familiar examples: phones, personal computers (PCs), [^7] Throughout the paper, we use “PC” to refer personal computers. We do not use the narrower definition of PC as a personal computer compatible with the original IBM PC. In particular, we consider Macs to be PCs. and video games. One clear strategy—arguably the only foolproof one—that would achieve the aim of protecting users from harms from third-party software would be to prevent phones from being able to download any software, and perhaps even bar them from the internet. Such an approach would, of course, undermine the functionality of the product and be a bad economic move. Existing smartphone providers strike different trade-offs, but none of them take an absolute approach.
@@
Whereas operating systems (OSs) for PCs allow users to download almost anything, video-game consoles have much more stringent control over what is allowed on the machine. This is not because of fundamental security distinctions between PCs and Xboxes, but rather because the respective developers of these technologies have chosen different trade-offs between security, interoperability, and other interests, taking into account the relevant application contexts.
PCs are key infrastructure for vast swathes of the economy and any exacting controls enacted by their manufacturers would likely either make the devices undesirable, if there is healthy competition, or negatively impact on the rest of the market, if there are monopoly conditions.
@@
In Section 3.2, we provide four case studies of security vetting concerns: app sideloading, in-app purchases, alternative app stores, and WebKit.
2.1.3 Hybrid Concerns
@@
Finally, there are situations featuring aspects of both security engineering concerns and security vetting concerns, in which interoperation would entail both the development of a secure system for interoperation and some sort of vetting of those developers looking to use that newly open system.
@@
These instances tend to feature a more complex relationship between the company and the third-party. In the security engineering setting, parties have to decide how to interoperate similar services and none of the parties in question necessarily has the upper hand in the decision making (although one company often has more power in practice). In the security vetting setting, the company that produces the platform is typically in power. The third-parties looking to provide their product on that platform must convince the company that their product is secure. We discuss these relationships further, from an economic perspective, in Section 2.2.
@@
Template of a Hybrid Concern From a platform’s perspective 1. Interoperating would mean allowing third parties to access functionality reserved for only our use. 2. Opening up those functionalities entails third parties newly interacting with security-sensitive parts of our OS. Designing a new system to secure such access is a costly engineering task that may require innovation. 3. Therefore, third-party access to these functionalities carries inherent security risk, and could undermine the security of billions of users.
@@
The hybrid category combines both of these aspects. The third-parties must still demonstrate that their products meet the requirements of an interoperation policy, but there is also more engineering and collaboration needed to support technical interoperability. For example, in opening mobile wallet functionality to third parties, device/OS providers need to both add APIs to allow developers access to previously restricted hardware and vet which developers will have access.
Section 3.3 provides two case studies of hybrid concerns: NFC for mobile payments and physical device interoperability.
2.1.4 Summary
@@
Our Templates are designed to capture just the essential elements of the argument patterns we have identified. Of course, they are oversimplifications with respect to concerns being raised in any real-world “security vs. interoperability” context. Unsurprisingly, the arguments made in practice are more complex, and may not neatly fall into a single category.
@@ 2.2 Horizontal and Vertical Interoperation
Our distinction between engineering and vetting concerns is closely correlated to distinctions in the relationships of the interoperating products.
@@
Horizontal interoperation happens when similar products need to be made interoperable. Messaging is a good example, where previously competing products must collaborate on integrated functionality and security (Case Study 3.1). This happens when the products are substitutable, a term used in economics to mean two products used for similar purposes that compete for the same user demand.
Vertical interoperation happens between products on different layers. Instances of vertical interoperation that arise naturally, as opposed to through regulatory or antitrust enforcement, tend to be between some platform and a complementary product. Again using economic terminology, we say two products are complementary if they are often used together, or, formally, if a price increase of one decreases the demand for the other. Apps are a particularly good example, and the success of smartphones is greatly dependent on the wide availability of apps and other complementary products that are developed by third parties.
Unlike with horizontal interoperability, there is an incentive to achieve good vertical interoperation between platforms and the apps that run on them, as these two products complement each other and increase customer utility in concert. Rather, the competitive friction to vertical interoperation is either from the platform wanting to hold onto extractionary practices (e.g., by taking commission on in-app purchases; see §3.2) or from an effort to protect the platform’s own complementary, higher layer product that is somehow privileged on the platform, which is called self-preferencing (e.g., payment apps having unique access to hardware functionality; see §3.3). This second case, where there is an in-house, higher layer, complementary product with exclusive access to platform functionality, is a hybrid case; the platform and the third-party product are complements, but the company’s higher layer product is a substitute.
@@
Relation to taxonomy
The economic context of horizontal or vertical interoperation may be useful to inform contextual analysis of “security vs. interoperability” arguments. Security engineering arguments are often related to horizontal interoperation, and security vetting arguments are often related to vertical interoperation, although this is not always the case. [^9] Theoretically, both types of arguments can arise in both horizontal and vertical interoperation scenarios. In practice, many complex real-world scenarios exhibit aspects of both, but the observation we wish to highlight is that there is a stronger correlation between certain categories.
Beyond our framework
We emphasize that the tension between security and interoperability is not inherent or universal. Plenty of technologies have been designed to support open interoperation from the ground up, and many the concerns we raised in this section will not arise at all. Interoperation between email providers is a good (horizontal) example; third-party software on Linux is a good (vertical) example illustrating that interoperation and security can go hand-in-hand.
@@ 3 Case Studies and Analysis
In this section, we examine examples from recent antitrust proceedings and discuss their details and how they fit into our framework. The proceedings are at different stages; some are complete while others are just getting started. Our focus is not on the details of the legal proceedings, but rather on the “security vs. interoperability” considerations in each case.
For each category in our taxonomy—engineering (§3.1), vetting (§3.2), and hybrid (§3.3)—we provide case studies and an analysis of patterns across the case studies, structured around four critical questions: (1) How do interoperation and economic incentives clash (or not)? (2) How do security arguments and economic incentives align (or not)? (3) What needs to change in order for the company to support interoperation? (4) What tension exists between security and other goals both before and after?
@@ 3.1 Security Engineering Concerns
This category is, in some sense, the simplest way to think about interoperation; here we have two or more parties trying to build or retrofit their systems to work together. Picture dual-gauge trains with different wheel gauges able to operate (safely) on tracks of different width. In our contexts, companies involved are generally competitors and the interoperation is generally horizontal.
This section, unlike the others, contains only one case study. Our case study contains two examples, iMessage and WhatsApp, which involve distinct considerations. We believe interoperable encrypted messaging is currently the most illustrative example in the category of security engineering concerns.
@@
There have been efforts to mandate interoperation between end-to-end encrypted (E2EE) messaging services in both the US and the EU, with a focus on iMessage (Apple) and WhatsApp (Meta), respectively. [^10] We focus on: United States v. Apple, Inc., No. 2:24-cv-04055 (D.N.J. filed Mar. 21, 2024) and CASE (EU) DMA.100024 Meta – number-independent interpersonal communications services, 2023 Interoperation would mean that users of one service could send an end-to-end encrypted (E2EE) message to the users of a different E2EE service, while preserving security [^11] REGULATION (EU) 2022/1925 on contestable and fair markets in the digital sector (Digital Markets Act), Article 7(3) .
Both WhatsApp and iMessage provide E2EE text messaging as a core part of their functionality. Like most modern messaging applications, they also offer other functionality: e.g., voice messages, voice calls, groups, and so on.
Before the actions of the antitrust authorities, WhatsApp users could only message other users on WhatsApp. In contrast, iMessage users could already send and receive messages originating from Android phones; however this type of messaging is not end-to-end encrypted (unlike iMessage chats between iPhone users) and generally has less functionality. The changes in response to the antitrust action are still materializing, but have mostly yet to be rolled out, so this reflects the status quo.
@@
Even when narrowing our focus to just E2EE text messaging, it is clear there is substantial engineering work to be done to achieve secure interoperation. Meta based WhatsApp encryption on the Signal Protocol, which is developed by the non-profit Signal Foundation, and is considered the gold standard in the field [51]. Meta will require that any third-party looking to interoperate uses the Signal Protocol or ‘a compatible protocol if they are able to demonstrate it offers the same security guarantees’ [51]. For many smaller providers, the possibility of interoperating with WhatsApp may provide ample incentive to do the engineering work to meet these requirements. For WhatsApp, secure interoperation will also require some engineering, including building secure connections with third-party servers and coordinating message formats [51]. This may take significant effort, but appears very tractable.
@@
Conversely, iMessage does not use the Signal protocol, and is sufficiently dominant in the market that it is not incentivized to either switch to the Signal protocol or otherwise facilitate interoperation. Historically, protocols for communication have been agreed upon by whole communities of users and operators, as in the case of email. The Signal Protocol is an effort in that direction and we might eventually have a system of community standards for secure messaging that everyone adheres to and that allows anyone to participate. Unlike email, messaging needs to be ‘retrofitted’ to allow for interoperation. For now, the efforts for greater interoperation consist of a combination of companies retrofitting for interoperation on the condition of adherence to their standards, and community-led protocol development efforts. There are real technical and research challenges, and it seems possible that the community can develop much better ways of achieving interoperation securely [43, 23].
@@
For the incumbents, interoperation entails allowing their services to send and receive information outside of their own networks, and thereby potentially helping competitors. These incumbent services are currently sufficiently popular that not allowing greater interoperation has been a boon – encouraging greater uptake – rather than a deterrent for customers. With the example of Apple and iMessage in particular, the incompatibility of messaging may play a role in users switching to iPhones because iMessage is only available on iPhones [35].
@@
One prominent point in the iMessage antitrust case [^12] United States v. Apple (see footnote 10). has been the difference in appearance between messages sent between iPhones (blue bubbles) and those sent from other devices (green bubbles) [35] and the associated impact on consumer choice. This issue is not directly related to the technical feasibility of interoperation at all. In practice, the color differentiation signals a lack of interoperation, and that green messages are not end-to-end-encrypted [1, 18].

### v2

@@ 3 Our Taxonomy
Security protections of existing
@@
TABLE II: Some Key Differences Between Security Engineering Concerns and Security Vetting Concerns
We now introduce our three-part taxonomy, which distinguishes two broad types of “security vs. interoperability” arguments as raising either security engineering concerns or security vetting concerns, summarized in Table II, alongside a third hybrid category that raises concerns incorporating aspects of both other categories. Our taxonomy clarifies the different types of security concerns that may arise in response to the prospect of introducing interoperation between products or systems which did not previously interoperate.
The technical, economic, and policy contexts in which these concerns are raised is critical to our taxonomy and subsequent analysis (§§45). In particular, the status quo of how the systems in question work before interoperation is introduced, the nature of the technical security guarantees the systems already provide, and the security policies and tradeoffs already in place, are essential considerations.
Security Engineering Concerns
When considering interoperating systems that are currently not technically interoperable, there are contexts in which preserving security alongside interoperation requires that some aspect of the company’s product or architecture be non-trivially changed. How to change systems to preserve the security guarantees currently offered, while also supporting interoperation, may or may not be clear based on existing techniques.
@@
Template of a Security Engineering Concern 1. Interoperating would require a building new technical mechanism or architecture for our system to interface with other systems. 2. Our system has existing technical security guarantees. Designing a new architecture to preserve these guarantees is an engineering task that is costly in time, effort, and coordination with others—and may require innovation, so success is not certain. 3. Therefore, interoperating secure systems is more costly than interoperating systems in general, and if we don’t get the engineering right, the security of billions of users could be undermined.
Security Vetting Concerns
Security vetting concerns arise in contexts where there is already some kind of interoperation supported by a system’s technical architecture. Many devices and platforms support an ecosystem of apps, software, and connected devices that are integral parts of their functionality, but that are developed by third parties. Examples include platforms, browsers, or operating systems on which third-party software can run, and platforms that offer API access to third-party developers.
Given such an architecture, there remains the question of which third parties a platform is willing to interoperate with. Downloads and other connections can open a vector of attack for malware, spyware, and other undesirable outcomes for both the user and the device. Companies may vet and control which apps, systems, or software are permitted to interact with their devices with a view to preventing or mitigating such undesirable security outcomes—although these decisions might be motivated by other goals too.
@@
Template of a Security Vetting Concern 1. Interoperating would mean our users and systems interacting with third parties and third-party software. 2. Third parties and third-party software may make different decisions about security than us. We cannot guarantee that third-party software would adhere to the same standards as software we have developed or carefully examined ourselves. 3. Therefore, allowing third-party software on our platform could undermine billions of users’ security.
@@
The above template is so generic that it could apply to any interoperation. As such, we find that simplistic vetting arguments may boil down to arguments directly against interoperation; in practice, the detailed context in which a vetting concern is raised is critical to a thorough assessment.
@@
Vetting third parties can be an important way to protect users, but also concentrates power and discretion of incumbent companies over ecosystems they control. Companies in this position can use their role as security guarantor to charge developers for access or prevent competitors from reaching the market. Platforms’ interoperation policy choices thus require striking a tradeoff between security, interoperation, and other business considerations.
@@
Due to the discretionary tradeoffs involved, arguments that security considerations necessitate a particular form of vetting or security measures generally cannot be substantiated by purely technical reasoning. Rather, they must refer to existing interoperation policies and the discretionary choices already present therein. This contrasts with security engineering concerns, which technical reasoning may suffice to substantiate, at least in some cases.
@@
Let us concretize the last point with a few familiar examples: phones, personal computers (PCs), [^12] Throughout the paper, we use “PC” to refer personal computers. We do not use the narrower definition of PC as a personal computer compatible with the original IBM PC. In particular, we consider Macs to be PCs. and video games. One clear strategy—arguably the only foolproof one—that would achieve the aim of protecting users from harms from third-party software would be to prevent phones from being able to download any software, and perhaps even bar them from the internet. Such an approach would, of course, undermine the functionality of the product and be a bad economic move. Existing smartphone providers strike different tradeoffs; none take an absolute approach.
@@
Whereas PC operating systems (OSs) allow users to download almost anything, video-game consoles have much more stringent control over what is allowed on the machine. This is not because of fundamental security differences between PCs and Xboxes, but rather because the respective developers of these technologies have chosen different tradeoffs between security, interoperability, and other interests, taking into account the relevant application contexts.
PCs are key infrastructure for vast swaths of the economy and valued for versatile multi-purpose use. Exacting controls enacted by their manufacturers could make the devices undesirable and therefore not competitive in healthy market conditions, or, under monopoly conditions, deeply harm both users and the software industry. This is all the more so given historically permissive norms around what software can run on PCs.
@@
In Section 4.2, we provide three case studies of security vetting concerns: app sideloading, alternative app stores, and browser engines (WebKit).
Hybrid Concerns
@@
Finally, there are situations featuring aspects of both engineering concerns and vetting concerns, in which interoperation entails both the development of a secure system for interoperation and some sort of vetting of those developers looking to use that newly open system.
@@
These instances tend to feature a more complex relationship between the company and the third party. In the security engineering setting, parties have to decide how to interoperate similar services, and none of the parties involved necessarily has the upper hand in decision making (although one company often has more power in practice). In the security vetting setting, the platform is typically in power. The third parties looking to offer a product on that platform must convince the company behind the platform that their product is secure. We discuss these relationships further, from an economic perspective, in Section 3.1.
@@
Template of a Hybrid Concern 1. Interoperating would mean allowing third parties to access functionality reserved for only our use. 2. Opening up those functionalities means third parties newly interacting with security-sensitive parts of our OS. Designing new systems to secure such access is a costly engineering task and may require innovation. 3. Therefore, third-party access to these functionalities carries inherent security risk, and could undermine the security of billions of users.
@@
The hybrid category combines both of these aspects. Third parties must still demonstrate that their products meet the requirements of an interoperation policy, but there is also more engineering and collaboration needed to support technical interoperability. For example, in opening mobile wallet functionality to third parties, device/OS providers need to both add APIs to allow developers access to previously restricted hardware and vet which developers will have access. Section 4.3 offers two case studies of hybrid concerns: NFC for payments and physical device interoperability.
Summary
@@
Our Templates are designed to capture just the essential elements of the argument patterns we have identified. Of course, they are oversimplifications with respect to concerns being raised in any real-world context. Unsurprisingly, the arguments made in practice are more complex, and may not neatly fall into a single category.
@@ 3.1 Horizontal and Vertical Interoperation
Our taxonomy is closely correlated to distinctions in the relationships of the interoperating products—horizontal and vertical relationships—corresponding to similar terminology that often appears in economics and antitrust law.
@@
Horizontal interoperation happens when similar products need to be made interoperable. Messaging is a good example, where previously competing products must collaborate on integrated functionality and security (Case Study 4.1). This happens when the products are substitutable, a term used in economics to mean two products used for similar purposes that compete for the same user demand [57]. That does not necessarily mean that the products are completely the same – one cannot yet use one messaging app to text users of a different messaging app. Formally, two goods are substitutable if the demand for one increases when the price for another increases. For example, if one messaging app introduces a paywall, we could imagine user demand shifting to a similar service.
Vertical interoperation happens between products on different layers. Instances of vertical interoperation that arise naturally (as opposed to through regulatory or antitrust enforcement) tend to be between some platform and a complementary product. Again using economic terminology, we say two products are complementary if they are often used together, or, formally, if a price increase of one decreases the demand for the other[57]. Smartphone apps are a good example, and the success of smartphones is greatly dependent on the wide availability of apps and other complementary products that are developed by third parties.
There is a clear incentive to achieve good vertical interoperation between platforms and the apps that run on them, as these products complement each other and increase customer utility in concert [63]. (Incentives in horizontal interoperation can be less cooperative.) Competitive friction in vertical interoperation may arise either from a platform maintaining extractionary practices (e.g., by taking commission on in-app purchases; see §4.2) or from a platform’s efforts to protect its own complementary, higher-layer product that is somehow privileged on its platform—this is called self-preferencing (e.g., payment apps having unique access to hardware functionality; see §4.3). The case where there is an in-house, higher-layer, complementary product with exclusive access to platform functionality, is a hybrid case; the platform and the third-party product are complements, and the company’s higher-layer product is a substitute.
Relation to taxonomy
The economic context of horizontal or vertical interoperation may be useful to inform contextual analysis of SvI arguments. Security engineering arguments often relate to horizontal interoperation, and security vetting arguments are often, though not always, related to vertical interoperation. [^14] Theoretically, both types of arguments can arise in both horizontal and vertical interoperation scenarios. In practice, many complex real-world scenarios exhibit aspects of both, but the observation we wish to highlight is that there is a stronger correlation between certain categories. Figure 1 illustrates these correlations.
@@
Scope of our taxonomy
Tensions between security and interoperability are not inherent or universal. Our work focuses on security concerns in the context of adding interoperability to systems that previously had less or no support for interoperation. Plenty of secure technologies have been designed to support open interoperation from the ground up, and many of the types of concerns discussed in this section will not arise at all in such cases. Interoperation between email providers is a good (horizontal) example; third-party software on Linux is a good (vertical) example illustrating that interoperation and security can go hand-in-hand.
@@ 4 Case Studies and Analysis
Based on recurring themes from the SvI discourse we reviewed, we propose four critical questions comprising an analytical framework for SvI concerns.
  1. 1.
    How do interoperation and economic incentives clash (or not)?
  2. 2.
    How do security arguments and economic incentives align (or not)?
  3. 3.
    What needs to change in order for the company to support interoperation?
  4. 4.
    What tension exists between security and other goals both before and after?
Next, for each category in our taxonomy—engineering (§4.1), vetting (§4.2), and hybrid (§4.3)—we provide one to three selected case studies, and analyze each case study with the four preceding questions in mind. Then in Section 5, we transition from individual case analysis to comparative analysis, structured around the same four questions.
@@ 4.1 Security Engineering Concerns
In security engineering cases, we have parties trying to build or retrofit their systems to work together. Picture dual-gauge trains with different wheel gauges being engineered to be able to operate (safely) on tracks of different width. In our contexts, the companies involved are generally competitors and the interoperation is generally horizontal.
@@
This case study discusses two examples of E2EE messaging, iMessage and WhatsApp, which involve distinct SvI considerations. Here, interoperation means that users of one service could send an end-to-end encrypted (E2EE) message to the users of a different E2EE service—while preserving security [^15] REGULATION (EU) 2022/1925 on contestable and fair markets in the digital sector (Digital Markets Act), Article 7(3) . We encountered mandated interoperability of Messenger (Meta) directly in our case search. There are also similar mandates for WhatsApp (Meta) in the EU and iMessage in the US [^16] CASE (EU) DMA.100024 Meta – number-independent interpersonal communications services, 2023 and United States v. Apple, Inc., No. 2:24-cv-04055 (D.N.J. filed Mar. 21, 2024). These three cases have many similarities so we discuss them together. We found SvI arguments stating that the interoperation is technically difficult and that third parties might not provide adequate security guarantees [61, 41].
Before the actions of the antitrust authorities, WhatsApp and Messenger users could only message other users on WhatsApp and Messenger, respectively. WhatsApp has since taken limited measures to support interoperation; Messenger remains siloed. iMessage is only available on iPhones, but iMessage users could always send and receive (SMS) messages to and from Android phones.171717iMessage-to-iMessage communication offers end-to-end encryption and additional features not available in iPhone-to-Android communication on iMessage. Unlike iMessage, WhatsApp and Messenger do not support SMS messaging. The three apps are used for similar purposes in practice.
@@
There is substantial engineering work to be done to achieve secure interoperation across E2EE messaging applications—even for basic messaging functionality, putting aside common auxiliary features such as disappearing messages, read receipts, and so on. Meta’s encryption for both WhatsApp and Messenger is based on the Signal Protocol, which is developed by the non-profit Signal Foundation, and is considered the gold standard in the field [61]. Meta has stated a requirement that third parties seeking to interoperate must use the Signal Protocol or ‘a compatible protocol if they are able to demonstrate it offers the same security guarantees’ [61]. For smaller providers, the possibility of interoperating with WhatsApp may provide ample incentive to do the engineering work to meet these requirements. For Meta, secure interoperation will also require some engineering, including building secure connections with third-party servers and coordinating message formats [61].
@@
In contrast, iMessage does not use the Signal protocol [52], and has sufficient market dominance that it may not be incentivized to either switch to the Signal protocol or otherwise facilitate interoperation. Historically, many communication protocols have been agreed upon by broad communities of users and operators, as with email. The Signal protocol is an effort in that direction and we might eventually have a system of community standards for secure messaging. But unlike email, messaging needs to be ‘retrofitted’ to support interoperation. For now, the efforts for greater interoperation consist of some companies retrofitting interoperation conditional on adherence to their standards, and some community-led protocol development efforts. There are real technical and research challenges, and it seems possible that the community can develop much better ways of interoperating securely [49, 24].
@@
For the incumbents, interoperation entails allowing their services to send and receive information outside of their own networks, and thereby potentially helping competitors. Major incumbent services are currently sufficiently popular that not allowing greater interoperation has been a boon—encouraging greater uptake—rather than a deterrent for customers (e.g., [40]). With the example of Apple and iMessage in particular, the incompatibility of messaging may play a role in users switching to iPhones because iMessage is only available on iPhones [38].
@@
One prominent point in the iMessage antitrust case [^18] United States v. Apple (see footnote 16). has been the difference in appearance between messages sent between iPhones (blue bubbles) and those sent from other devices (green bubbles) [38] and the associated impact on consumer choice. This issue is not directly related to the technical feasibility of interoperation at all. In practice, the color differentiation signals a lack of interoperation, and that green messages are not end-to-end-encrypted [1, 18].